HUBBLE SPACE TELESCOPE DAILY REPORT #847 PERIOD COVERED: 0800 EST Daily Status Report as of 078/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 3755 (Absorption Cross-Section of Nearby Galaxies) The High Resolution Spectrograph was used to attempt to make observations of quasar 1543+489 which is to be used as a probe to study the absorption lines produced by the halos of a nearby galaxy. There were six losses of lock while the gyros were being run in the High Rate Mode. Lock was recovered each time. The observations was received and displayed in the OSS. It image was reported to show signs of spacecraft motion which resulted from the high gyro rate test. In spite of the problems there were a number of features visible in the spectra. 1.2 Completed Three Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies With FOC/96) The Faint Object Spectrograph was used to attempt to take ultraviolet snapshot images of two nearby galaxies UGC-6516 and ESO-1006-2. Both observations could not be made because the baseline guide star acquisitions failed. (See Not 2.1) 1.3 Completed GHRS/2 1177 (Search For Hot Plasmas in Outer Atmosphere) The High Resolution Spectrograph was used to make spectral observations looking for evidence of very hot (>10,000 Degs. K) plasma in the outer atmosphere of the giant K class star HD48329. The first series of observations were not made because the baseline guide star acquisition and subsequent reacquisitions failed. (See Note 2.1) The deflection calibration did not occur. The spiral search target acquisition was received in real time. The target was found, and the flux level was reported to be 95,487. The The next observations were successful. The wavelenght calibrations and the spectral observations were executed without problems. The observations were received and displayed. they were reported to look okay. 1.4 Uplinked Updated Ephemeris Tables, Per MOSES 2871 2.0 FLIGHT OPERATIONS SUMMARY: There was a series of acquisition, reacquisition, and Fixed Head Star Tracker (FHST) failures following a Type 2 slew which produced large attitude errors. 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 3 (Failures at 077/1543Z, 1705Z, and 1823Z) Scheduled Reacquisitions : 9 Successful : 7 (Failures at 077/2000Z, and 078/0404) Losses of Lock : 6 (Between 0200Z and 0220Z, due to High Rate Gyro Test. See Note 1.1 ) 2.2 FHST Updates: Scheduled : 10 Successful : 3 (Failures at 077/1537Z,1540Z 1644Z, 1647Z, 1738, 1800Z, and 1803Z) 2.3 Significant HSTARs: None 2.4 Operations Note: There were two brief reoccurrences of the Magnetometer #1 V2 channel saturation anomaly on Day 069. They lasted for 10- 30 seconds each. These incidents took place when the HST was at a sun angle of approximately 51 degrees. A geomagnetic analysis has shown that the MSS electronics was shadowed by the aperture door resulting in cold temperatures. This will occur when the sun angle is less than 57 degrees. The last portion of the executing 074 SMS was replanned to eliminate observations that were in this zone. Additionally, future SMSs are being reworked by the STScI to remove observations in this region. A high mode gyro bias calibration was performed on Day 075 for gyros 2,3 and 5. The data changes of (12, 5, 573) arcseconds per hour. This represents a higher value than over previous bias estimates. The bias data was not uplinked. Three more bias attempts, during course track control, failed to provide useful data. On Day 078, another attempt at calibrating the high mode bias was successful for gyros 3,4 and 5. The data indicates changes of (-2, 0, and 188) arcseconds per hour. The data from this HGBU was uplinked to the spacecraft. Two more high bias calibrations will be performed on Days 078 and 079. Scale factor estimates are also planned. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 3496 (Ultraviolet Atlas of Sirius) 3.2 Complete Two Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) 3.3 Complete PC 4769 (Structure of Quasars and Related Objects) 3.4 Complete WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) 3.5 Complete GHRS/2 3468 (UV Molecular Band in HD44179) 3.6 Complete GHRS/2 3608 (12c-13c Isotope Ratio in Planetary Nebulae) 3.7 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.8 Complete PC 3872 (Blue Stragglers in the Cores of Globular Clusters) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: FGS #2 Anomaly Investigation Further testing and analysis of the FGS #2 anomalous behavior are still ongoing. In the interim FGS #2 will not be used for observations. 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-